# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ 3.3. Eligibility Criteria
    • The study must focus on at least one of the following NLP tasks: Legal Question Answering, Legal Named Entity Recognition, Legal Judgment Prediction, Legal Document Summarization, Legal Text Classification, Legal Language Models, or legal corpora.
@@
    • Both peer-reviewed and non-peer-reviewed (e.g., arXiv) studies were considered if they provided valuable insights.
@@
    • Studies focused exclusively on unrelated areas such as information retrieval methods, pattern mining, information extraction, or similarity detection without a clear application to the specific legal NLP tasks mentioned.
@@
4. NLP tasks, datasets, and large corpora in legal domains
NLP tasks in the legal domains cover a range of specialized applications that address the unique challenges and requirements of legal texts. These tasks leverage advanced NLP techniques to process, analyze, and extract meaningful information from legal documents. Legal NLP tasks help make legal research and decision-making more efficient and accurate. In this section, we explore various NLP tasks that are tailored to the legal domain and show their impact on the legal section. Additionally, we will discuss existing works and research related to each task and provide an overview of the current state of the art in this field.
Furthermore, the success of these NLP tasks heavily depends on the availability and quality of domain-specific datasets. Therefore, we will also examine the datasets commonly used in legal NLP research, exploring their characteristics and the role they play in advancing the development of NLP models for the legal domain.
4.1. Legal Question Answering
4.1.1. Task
@@
LQA involves responding to queries about law. This task is typically done by legal professionals skilled in the relevant legal field. This process requires a comprehensive review of existing laws, careful interpretation of statutes and regulations, and the application of legal principles and past cases to the relevant facts. LQA seeks to offer precise advice on legal matters. It helps people and businesses in navigating the legal landscape and addressing legal challenges. Recently, DL has been leveraged in LQA to employ neural network models that train on extensive datasets to identify complex patterns and relationships. These models evaluate the questions posed, identify relevant legal topics, and produce suitable answers based on the patterns they have learned.
@@
Modern ML approaches for LQA, particularly DL, rely on neural networks to understand natural language. Popular architectures include Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Convolutional Neural Networks (CNN), which can be fine-tuned for specific tasks such as QA. These models generate accurate responses, adapt to new patterns, understand context, and provide relevant answers. Transformer-based models, such as BERT (Devlin et al., 2019) and ChatGPT (Ouyang et al., 2022), have proven particularly effective in NLP tasks. These models use the transformer architecture and self-attention mechanisms to learn the context of the text and understand the meaning of words. This allows them to provide relevant answers by effectively weighing the importance of different parts of the input. In the following, we will study the existing LQA works in legal domain.
@@
Huang et al. (2020) introduce the Artificial Intelligence Law Assistant, the first Chinese LQA system that integrates a legal knowledge graph (KG) to enhance query comprehension and answer ranking. The system collects a large-scale QA corpus from an online legal forum and constructs a legal KG with over 42,000 legal concepts. It employs a knowledge-enhanced interactive attention network using Bi-LSTM and co-attention mechanisms to enrich semantic representations of QA pairs with legal domain knowledge. Additionally, it provides visual explanations for selected answers, offering users a clear understanding of the QA process.
@@
An example of an answer retrieval system specific to Private International Law is proposed by Sovrano et al. (2020). This system integrates Term Frequency Inverse Document Frequency (TF-IDF) with deep LMs to retrieve relevant answers from an automatically generated KG of contextualized grammatical sub-trees. The KG aligns with a legal ontology based on Ontology Design Patterns, such as agent, role, event, temporal parameter, and action, to reflect the legal significance of the relationships within and between provisions.
@@
Khazaeli et al. (2021) develop an IR-based QA system tailored to the legal domain, combining sparse vector search (BM25) and dense vector techniques (semantic embeddings) as input to a BERT-based (Devlin et al., 2019) answer re-ranking system. The system utilizes Legal GloVe and Legal Siamese BERT embeddings to enhance retrieval effectiveness. An answer finder component computes the probability of a passage answering the question using a BERT sequence binary classifier fine-tuned on question-answer pairs, improving the model’s ability to discriminate good answers.
@@
Li et al. (2024b) introduce a retrieving-then-answering framework featuring the Graph-Based Evidence Retrieval and Aggregation Network (GESAN) to enhance QA on the Judicial Examination of Chinese Question Answering (JEC-QA) dataset (Zhong et al., 2020c). The framework leverages relevant legal knowledge by predicting question topics and retrieving legal paragraphs using BM25. GESAN aggregates the evidence and processes it along with the question and options to make accurate predictions, demonstrating improved accuracy and reasoning abilities in LQA.
@@
Askari et al. (2022) propose a method for generating query-dependent textual profiles for lawyers to improve legal expert finding on QA platforms. Using data from the Avvo [^1] https://www.avvo.com QA forum, they focus on aspects such as sentiment, comments, and recency to create profiles. These profiles are fine-tuned using BERT models (Devlin et al., 2019), and the final aggregated score is calculated using a linear combination of profile-trained BERT models. It improves retrieval performance in the legal expert finding task.
@@
Zhang et al. (2023b) propose a generation-based method for LQA, modeling it as a generation task to produce new, relevant answers tailored to each question. The system incorporates laws as external knowledge into the answer generation process, using a retriever to fetch applicable law articles and a generator to create answers using this knowledge. Both components are integrated into a single T5 (Raffel et al., 2020) model using MTL. It can enhance the model’s understanding and generation abilities while ensuring answers are accurate and informative.
@@
Sovrano et al. (2024) propose DiscoLQA, a discourse-based LQA system that focuses on important discourse elements like Elementary Discourse Units (EDUs) and Abstract Meaning Representations (AMRs). This approach helps the answer retriever identify the most relevant parts of the discourse, enhancing retrieval accuracy. They introduce the Q4EU dataset, containing over 70 questions and 200 answers on six European norms, demonstrating improved performance in legal QA even without domain-specific training.
Yuan et al. (2023) present a three-step approach to bridge the legal knowledge gap by creating CLIC-pages—snippets that explain technical legal concepts in layperson’s terms. They construct a legal question bank containing legal questions answered by CLIC-pages, using large-scale PLMs like GPT-3 (Brown et al., 2020) to generate machine-generated questions. The study demonstrates that machine-generated questions are more scalable and diversified, aiding in improving accessibility to legal information for non-experts.
Askari et al. (2024) propose a cross-encoder re-ranker ($CE_{FS}$) for legal answer retrieval, incorporating fine-grained structured inputs from community QA data to enhance retrieval effectiveness. They introduce the LegalQA dataset containing 9,846 questions and 33,670 lawyer-curated answers. The approach involves a two-stage ranking pipeline with a BM25 retriever followed by a re-ranker, showing that integrating question tags into the input structure can bridge the knowledge gap and improve retrieval in the legal domain.
4.1.2. Datasets
The LQA datasets are a specialized resource designed to facilitate research in the domain of legal NLP. they consist of a collection of legal questions and corresponding answers, drawn for various legal documents and case law. Most questions in the LQA datasets fall into two main categories: knowledge-driven questions (KD-questions) and case-analysis questions (CA-questions) (Zhong et al., 2020c). KD-questions are centered around the understanding of specific legal concepts, whereas CA-questions involve the analysis of actual legal cases. Both types demand advanced reasoning skills and a deep comprehension of the text, making LQA a particularly challenging task in the field of NLP.
Zhong et al. (2020c) introduce JEC-QA, a dataset with 26,365 multiple-choice questions from the National Judicial Examination of China and related websites. Each question offers four possible answers and is labeled with the type of reasoning required, such as word matching, concept understanding, numerical analysis, multi-paragraph reading, and multi-hop reasoning. This dataset poses significant challenges for QA models, highlighting the gap between machine performance and human expertise in complex legal reasoning.
@@ 4.1.2. Datasets
4.2. Legal Judgment Prediction
4.2.1. Task
@@
LJP is an important task within legal NLP, especially prevalent in civil law systems where judgments are determined based on case facts and statutory articles (Zhong et al., 2020b). This task involves predicting legal outcomes from the descriptions of cases and relevant legal statutes, and is essential in countries like France, Germany, Japan, and China (Zhong et al., 2020b). LJP has garnered significant attention from AI researchers and legal professionals due to its potential to assist judges, lawyers, and legal scholars in predicting case outcomes based on historical data (Chen et al., 2024a).
@@
Despite its promise, LJP is a highly demanding and challenging task. It requires careful handling of natural biases in historical legal data, which can create feedback loops that amplify discrimination (Krasadakis et al., 2024). Therefore, ensuring the impartiality of predicted rulings is crucial (Krasadakis et al., 2024). Currently, LJP is primarily performed by legal experts who undergo extensive specialized training to manage the complex steps involved, such as identifying relevant law articles, defining charge ranges, and deciding penalty terms (Cui et al., 2023). Nevertheless, LJP provides substantial benefits, streamlining legal decision-making processes for both practitioners and ordinary citizens (Cui et al., 2023).
@@
Luo et al. (2017) propose an attention-based neural network to enhance charge prediction by jointly modeling charge prediction and relevant law article extraction. They used Bidirectional Gated Recurrent Units (Bi-GRUs) to encode fact descriptions and an article extractor to identify top relevant law articles. The model employs an attention mechanism guided by context vectors to combine embeddings for prediction. Evaluations on Chinese judgment documents showed improved accuracy in predicting charges and providing relevant legal articles.
@@
Zhong et al. (2018) introduce TopJudge, a topological MTL framework that models dependencies among subtasks in LJP, such as law article prediction, charge prediction, and penalty terms. Using a Directed Acyclic Graph (DAG), TopJudge processes subtasks in a topological order reflecting real-world legal decision-making. Evaluated on large-scale Chinese criminal case datasets, it outperformed previous models in predicting legal outcomes.
@@
Ye et al. (2018) address the problem of Court View Generation from fact descriptions in criminal cases to enhance the interpretability of charge prediction systems and aid in automatic legal document generation. They formulated this as a text-to-text Natural Language Generation (NLG) problem, using a label-conditioned Seq2Seq model with attention to decode court views based on encoded charge labels. Their model outperformed basic Seq2Seq models in generating accurate and natural court views. This work contributes to automatic legal document generation by providing justifications for charge decisions.
@@
Yang et al. (2019) propose a Multi-Perspective Bi-Feedback Network (MPBFN) with a Word Collocation Attention mechanism to improve LJP. MPBFN addresses the challenges of multiple subtasks and their dependencies by using a bi-feedback mechanism for forward prediction and backward verification among subtasks. The Word Collocation Attention integrates word collocation features and numerical semantics to better predict penalties. Evaluated on the CAIL-small and CAIL-big datasets (Xiao et al., 2018), their model outperformed baselines in predicting law articles, charges, and penalty terms.
@@
Chalkidis et al. (2019a) introduce an English LJP dataset containing approximately 11.5k cases from the European Court of Human Rights (ECHR). They evaluated various neural models on this dataset, including a hierarchical version of BERT  (Devlin et al., 2019) (HIER-BERT) to handle long legal documents. Their models outperformed previous feature-based approaches in tasks like violation classification and case importance prediction. They also explored potential biases in legal predictive models using data anonymization.
@@
Medvedeva et al. (2020) investigate using NLP tools to predict judicial decisions of the ECHR based on court proceeding texts. They employed an SVM linear classifier to predict violations of articles, achieving an average accuracy of 75%. However, when predicting future decisions based on past cases, accuracy decreased to 58–68%. The study also found that predicting outcomes based solely on judges’ surnames could achieve an average accuracy of 65%, highlighting potential biases.
@@
Zhong et al. (2020a) introduce QAjudge, a reinforcement learning-based model designed to provide interpretable legal judgments by visualizing the prediction process. QAjudge uses a Question Net to iteratively select relevant yes-no questions about case facts, an Answer Net to provide answers, and a Predict Net to generate the final judgment. The model aims to minimize the number of questions asked. It focuses on crucial elements to ensure fairness and interpretability. Evaluated on real-world datasets, QAjudge demonstrated potential in providing reliable and transparent legal judgments.
@@
Xu et al. (2020) propose the Law Article Distillation based Attention Network (LADAN), an end-to-end model addressing the issue of confusing charges in LJP by distinguishing similar law articles. The model uses a novel graph neural network to learn differences between confusing law articles and an attention mechanism to extract discriminative features from fact descriptions. Experiments on real-world datasets showed that LADAN improved performance over previous methods in law article prediction, charge prediction, and penalty term prediction.
Strickson and De La Iglesia (2020) present the first LJP model for UK court cases, creating a labeled dataset of UK court judgments spanning 100 years. They evaluated various ML models and feature representations, with their best model achieving an accuracy of 69%. The study demonstrated the potential of LJP for UK courts, though challenges remain due to the complexity of legal language and lack of structured public datasets.
@@
Ma et al. (2021) introduce MSJudge, a MTL framework designed to predict legal judgments by leveraging multi-stage judicial data, including pre-trial claims and court debates. MSJudge consists of components to encode multi-stage context, model interactions among claims, facts, and debates, and predict judgments. Evaluated on a large civil trial dataset [^2] https://github.com/mly-nlp/LJP-MSJudge, MSJudge outperformed state-of-the-art baselines, enhancing trial efficiency and judgment quality.
Almuslim and Inkpen (2022) focus on LJP for Canadian appeal court cases, employing various NLP and ML methods to predict binary outcomes (‘Allow’ or ‘Dismiss’) based on case descriptions. Deep learning models using custom Word2Vec embeddings achieved the highest accuracy of 93%, significantly outperforming classical ML models. The study highlights the potential of predictive models to aid legal professionals and establishes a foundation for future research in the Canadian legal system.
@@
Feng et al. (2022) address limitations of state-of-the-art LJP models by proposing an event-based prediction model with constraints to improve performance. The model extracts fine-grained key events from case facts and predicts judgments based on these events rather than the entire fact statement. They manually annotated a legal event dataset and introduced output constraints to guide learning. Their method effectively leverages event information and cross-task consistency constraints, enhancing LJP accuracy.
@@
Previous works mainly focus on creating accurate representations of a case’s fact description to enhance judgment prediction performance. However, these methods often overlook the practical judicial process, where human judges compare similar law articles or potential charges before making a decision. To address this gap, Zhang et al. (2023a) propose CL4LJP, a supervised contrastive learning framework to improve LJP by capturing fine-grained differences between similar law articles and charges. The framework includes contrastive learning tasks at the article, charge, and label levels, enhancing the model’s ability to model relationships between fact descriptions and labels. Experiments demonstrated that CL4LJP outperformed previous methods, proving its effectiveness and robustness.
@@
4.2.2. Datasets
The LJP datasets are specialized resources designed to advance research in predicting judicial outcomes within the domain of legal NLP. These datasets are categorized into four main types: court view generation datasets, law articles, charge prediction, and prison term prediction. Court view generation datasets involve court opinions and summaries. Law Articles datasets involve the prediction of legal outcomes based on specific statutes or regulations. Charge prediction datasets are concerned with predicting the charges that should be brought against a defendant based on the case details, while prison term prediction datasets aim to estimate the likely sentence duration given the nature of the crime and the legal context. Each type of dataset presents unique challenges, demanding not only text comprehension but also the ability to apply complex legal reasoning, making LJP a particularly complex task in the field of NLP.

### v2

@@ 4.3. Eligibility Criteria
    • The study must focus on at least one of the following NLP tasks: LQA, legal NER, LJP, LDS, LTC, LAM, legal LMs, or legal corpora.
@@
    • Both peer-reviewed and non-peer-reviewed studies were considered if they provided valuable insights.
@@
    • Studies focused exclusively on unrelated areas such as IR methods, pattern mining, information extraction, or similarity detection without a clear application to the specific legal NLP tasks mentioned.
@@
5. Legal Question Answering
@@
LQA involves responding to queries about law. This task is typically done by legal professionals skilled in the relevant legal field. This process requires a comprehensive review of existing laws, careful interpretation of statutes and regulations, and the application of legal principles and past cases to the relevant facts. LQA aims to provide advice on legal matters. It helps people and businesses in navigating the legal landscape and addressing legal challenges.
5.1. Task
Recently, DL has been leveraged in LQA to employ NN models that train on extensive datasets to identify complex patterns and relationships. These models evaluate the questions posed, identify relevant legal topics, and produce suitable answers based on the patterns they have learned.
@@ 5.1. Task
Modern ML approaches for LQA rely on NN to understand natural language. Popular architectures include Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Convolutional Neural Networks (CNN), which can be fine-tuned for specific tasks such as QA. These models generate better responses, adapt to new patterns, understand context, and provide relevant answers. Transformer-based models, such as BERT and ChatGPT, have proven particularly effective in NLP tasks. These models use the transformer architecture and self-attention mechanisms to learn the context of the text and understand the meaning of words. This allows them to provide relevant answers by weighing the importance of different parts of the input. In the following, we will study the existing LQA works in legal domain.
@@
Huang et al. (2020) introduce the Artificial Intelligence Law Assistant, the first Chinese LQA system that integrates a legal Knowledge Graph (KG) to enhance query comprehension and answer ranking. The system collects a large-scale QA corpus from an online legal forum and constructs a legal KG with over 42,000 legal concepts. It employs a knowledge-enhanced interactive attention network using Bidirectional LSTM (Bi-LSTM) and co-attention mechanisms to enrich semantic representations of QA pairs with legal domain knowledge. Additionally, it provides visual explanations for selected answers, offering users a clear understanding of the QA process.
@@
An example of an answer retrieval system specific to Private International Law is proposed by Sovrano et al. (2020). This system integrates Term Frequency Inverse Document Frequency (TF-IDF) with deep LMs to retrieve relevant answers from an automatically generated KG of contextualised grammatical sub-trees. The KG aligns with a legal ontology based on Ontology Design Patterns, such as agent, role, event, temporal parameter, and action, to capture the relationships within and between provisions.
@@
Khazaeli et al. (2021) develop an IR-based QA system tailored to the legal domain, combining sparse vector search (BM25) and dense vector techniques (semantic embeddings) as input to a BERT-based answer re-ranking system. The system utilises Legal GloVe and Legal Siamese BERT embeddings to enhance retrieval performance. An answer finder component computes the probability of a passage answering the question using a BERT sequence binary classifier fine-tuned on question-answer pairs, improving the model’s ability to discriminate good answers.
@@
Li et al. (2024b) introduce a retrieving-then-answering framework featuring the Graph-Based Evidence Retrieval and Aggregation Network (GESAN) to enhance QA on the Judicial Examination of Chinese Question Answering (JEC-QA) dataset (Zhong et al., 2020c). The framework leverages relevant legal knowledge by predicting question topics and retrieving legal paragraphs using BM25. GESAN aggregates the evidence and processes it along with the question and options to make accurate predictions, demonstrating improved reasoning abilities in LQA.
@@
Askari et al. (2022) propose a method for generating query-dependent textual profiles for lawyers to improve legal expert finding on QA platforms. Using data from the Avvo [^2] https://www.avvo.com QA forum, they focus on aspects such as sentiment, comments, and recency to create profiles. These profiles are fine-tuned using BERT, and the final aggregated score is calculated using a linear combination of profile-trained BERT. It improves retrieval performance in the legal expert finding task.
@@
Zhang et al. (2023c) propose a generation-based method for LQA, modelling it as a generation task to produce new, relevant answers tailored to each question. The system incorporates laws as external knowledge into the answer generation process, using a retriever to fetch applicable law articles and a generator to create answers using this knowledge. Both components are integrated into a single T5 (Raffel et al., 2020) model using MTL. It can improve the model’s understanding and generation abilities while ensuring answers are accurate and informative.
@@
Sovrano et al. (2024) propose DiscoLQA, a discourse-based LQA system that focuses on important discourse elements, such as Elementary Discourse Units and Abstract Meaning Representations. This approach helps the answer retriever identify the most relevant parts of the discourse, enhancing retrieval accuracy. They introduce the Q4EU dataset, containing over 70 questions and 200 answers on six European norms, demonstrating improved performance in legal QA even without domain-specific training.
Yuan et al. (2023) present a three-step approach to bridge the legal knowledge gap by creating CLIC-pages—snippets that explain technical legal concepts in layperson’s terms. They construct a legal question bank containing legal questions answered by CLIC-pages, using GPT-3 (Brown et al., 2020) to generate machine-generated questions. The study shows that machine-generated questions improve scalability and access to legal information for non-experts.
Askari et al. (2024) propose a cross-encoder re-ranker ($CE_{FS}$) for legal answer retrieval, incorporating fine-grained structured inputs from community QA data to enhance retrieval performance. They introduce the LegalQA dataset containing 9,846 questions and 33,670 lawyer-curated answers. The approach involves a two-stage ranking pipeline with a BM25 retriever followed by a re-ranker, showing that integrating question tags into the input structure can bridge the knowledge gap and improve retrieval in the legal domain.
5.2. Datasets
The LQA datasets are a specialised resource designed to facilitate research in the domain of legal NLP. they consist of a collection of legal questions and corresponding answers, drawn for various legal documents and case law. Most questions in the LQA datasets fall into two main categories: knowledge-driven questions (KD-questions) and case-analysis questions (CA-questions) (Zhong et al., 2020c). KD-questions are centred around the understanding of specific legal concepts, whereas CA-questions involve the analysis of actual legal cases. Both types demand advanced reasoning skills and a deep comprehension of the text, making LQA a particularly challenging task in the field of NLP.
Zhong et al. (2020c) introduce JEC-QA, a dataset with 26,365 multiple-choice questions from the National Judicial Examination of China and related websites. Each question offers four possible answers and is labelled with the type of reasoning required, such as word matching, concept understanding, numerical analysis, multi-paragraph reading, and multi-hop reasoning. This dataset poses challenges for QA models, highlighting the gap between machine performance and human expertise in legal reasoning.
@@ 5.2. Datasets
6. Legal Judgement Prediction
@@
LJP is an important task within legal NLP, especially prevalent in civil law systems where judgements are determined based on case facts and statutory articles (Zhong et al., 2020b). This task involves predicting legal outcomes from the descriptions of cases and relevant legal statutes, and is essential in countries, such as France, Germany, Japan, and China (Zhong et al., 2020b). LJP has attracted interest from AI researchers and legal professionals due to its potential to assist judges, lawyers, and legal scholars in predicting case outcomes based on historical data (Chen et al., 2024a).
@@
Despite its promise, LJP is a highly demanding and challenging task. It requires careful handling of natural biases in historical legal data, which can create feedback loops that amplify discrimination (Krasadakis et al., 2024). Therefore, ensuring the impartiality of predicted rulings is crucial (Krasadakis et al., 2024). Currently, LJP is primarily performed by legal experts who undergo extensive specialised training to manage the complex steps involved, such as identifying relevant law articles, defining charge ranges, and deciding penalty terms (Cui et al., 2023). Nevertheless, LJP provides substantial benefits, streamlining legal decision-making processes for both practitioners and ordinary citizens (Cui et al., 2023).
6.1. Task
@@
Luo et al. (2017) propose an attention-based NN to enhance charge prediction by jointly modelling charge prediction and relevant law article extraction. They used Bidirectional Gated Recurrent Units (Bi-GRUs) to encode fact descriptions and an article extractor to identify top relevant law articles. The model employs an attention mechanism guided by context vectors to combine embeddings for prediction. Evaluations on Chinese judgement documents showed improved accuracy in predicting charges and providing relevant legal articles.
@@
Zhong et al. (2018) introduce TopJudge, a topological MTL framework that models dependencies among subtasks in LJP, such as law article prediction, charge prediction, and penalty terms. Using a Directed Acyclic Graph, TopJudge processes subtasks in a topological order reflecting real-world legal decision-making. Evaluated on large-scale Chinese criminal case datasets, it outperformed previous models in predicting legal outcomes.
@@
Ye et al. (2018) address the problem of Court View Generation from fact descriptions in criminal cases to enhance the interpretability of charge prediction systems and aid in automatic legal document generation. They formulated this as a text-to-text Natural Language Generation (NLG) problem, using a label-conditioned Seq2Seq model with attention to decode court views based on encoded charge labels. This work contributes to automatic legal document generation by providing justifications for charge decisions.
@@
Yang et al. (2019) propose a Multi-Perspective Bi-Feedback Network (MPBFN) with a Word Collocation Attention mechanism to improve LJP. MPBFN addresses the challenges of multiple subtasks and their dependencies by using a bi-feedback mechanism for forward prediction and backward verification among subtasks. The Word Collocation Attention integrates word collocation features and numerical semantics to better predict penalties. Evaluated on the CAIL datasets (Xiao et al., 2018), their model outperformed baselines in predicting law articles, charges, and penalty terms.
@@
Chalkidis et al. (2019a) introduce an English LJP dataset containing approximately 11.5k cases from the European Court of Human Rights (ECHR). They evaluated various neural models on this dataset, including a hierarchical version of BERT (HIER-BERT) to handle long legal documents. Their models outperformed previous feature-based approaches in tasks, such as violation classification and case importance prediction. They also explored potential biases in legal predictive models using data anonymization.
@@
Medvedeva et al. (2020) investigate using NLP tools to predict judicial decisions of the ECHR based on court proceeding texts. They employed Support Vector Machine (SVM) linear classifier to predict violations of articles, achieving an average accuracy of 75%. However, when predicting future decisions based on past cases, accuracy decreased to 58–68%. They also found that predicting outcomes based solely on judges’ surnames could achieve accuracy of 65%, highlighting potential biases.
@@
Zhong et al. (2020a) introduce QAjudge, a reinforcement learning-based model designed to provide interpretable legal judgements by visualising the prediction process. QAjudge uses a Question Net to iteratively select relevant yes-no questions about case facts, an Answer Net to provide answers, and a Predict Net to generate the final judgement. The model aims to minimise the number of questions asked. Evaluated on real-world datasets, QAjudge demonstrated potential in providing reliable and transparent legal judgements.
@@
Xu et al. (2020) propose the Law Article Distillation based Attention Network (LADAN), an end-to-end model addressing the issue of confusing charges in LJP by distinguishing similar law articles. The model uses a novel graph NN to learn differences between confusing law articles and an attention mechanism to extract discriminative features from fact descriptions. Experiments on real-world datasets showed that LADAN improved performance over previous methods in law article prediction, charge prediction, and penalty term prediction.
@@
Ma et al. (2021) introduce MSJudge, a MTL framework designed to predict legal judgements by leveraging multi-stage judicial data, including pre-trial claims and court debates. MSJudge consists of components to encode multi-stage context, model interactions among claims, facts, and debates, and predict judgements. Evaluated on a large civil trial dataset, MSJudge outperformed state-of-the-art baselines in judgement quality.
@@
Feng et al. (2022) address limitations of state-of-the-art LJP models by proposing an event-based prediction model with constraints to improve performance. The model extracts fine-grained key events from case facts and predicts judgements based on these events rather than the entire fact statement. They manually annotated a legal event dataset and introduced output constraints to guide learning. Their method leverages event information and cross-task consistency constraints.
@@
Previous works mainly focus on creating accurate representations of a case’s fact description to enhance judgement prediction performance. However, these methods often overlook the practical judicial process, where human judges compare similar law articles or potential charges before making a decision. To address this gap, Zhang et al. (2023a) propose CL4LJP, a supervised contrastive learning framework to improve LJP by capturing fine-grained differences between similar law articles and charges. The framework includes contrastive learning tasks at the article, charge, and label levels, enhancing the model’s ability to model relationships between fact descriptions and labels.
@@
6.2. Datasets
The LJP datasets are specialised resources designed to advance research in predicting judicial outcomes within the domain of legal NLP. These datasets are categorised into four main types: court view generation datasets, law articles, charge prediction, and prison term prediction. Court view generation datasets involve court opinions and summaries. Law articles datasets involve the prediction of legal outcomes based on specific statutes or regulations. Charge prediction datasets are concerned with predicting the charges that should be brought against a defendant based on the case details, while prison term prediction datasets aim to estimate the likely sentence duration given the nature of the crime and the legal context. Each type of dataset presents unique challenges, demanding not only text comprehension but also the ability to apply complex legal reasoning.